OpenStudy (anonymous):

please help!? What is the vertical asymptote of y=log(x-2)+5?

jimthompson5910 (jim_thompson5910):

You can't take the log of 0, so x-2 is positive (since you can't take the log of a negative either) if the argument of the log was zero, then... x - 2 = 0 x = 2 This value causes x-2 to be zero, which creates the vertical asymptote So the vertical asymptote is x = 2

OpenStudy (anonymous):

@jim_thompson5910 THANK YOU SO MUCH! you really helped me a lot :) can you help me with this one more just to be positive that i understand? i would really appreciate it! i can give you another best response? :) thanks so much anyway! What is the vertical asymptote of y=ln(x+3)+2?

jimthompson5910 (jim_thompson5910):

same idea, ln is a natural log though

jimthompson5910 (jim_thompson5910):

so x+3 = 0 solve for x and whatever you get is your vertical asymptote

OpenStudy (anonymous):

so x =-3?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

thanks so much again!! :)